1. Mr. Llanto buys a golden casing of iphone worth php 2,540 in dubai and tries to sell it in cagayan de oro. He planned to have a 60% markup based on cost then set the selling price based on this target. However, since no one is buying it. He marled down the casing to 20%. After another week, he marked it down for 10%. After a few weeks someone bought it.




A. What is the final selling price of the golden casing?




B. Given the final selling price for the casing, what is Mr. Llanto's markup rate based on cost?




2. Clarissa bought the following items: 250 grams of pckle relish at php. 0.40 per gram




1/2 kilogram of ham at php 275.00 per kilogram




200 grams of cheese at php 0.50 per gram




10 grams of bell pepper at php 135 per kilogram




6 pieces of eggs at php 7 per piece




If clarissa gave the cashier php 1,000, how much will be her change?




3. In school, there are 650 students. The ratio of the boys to the girls is 8:5. How many more girls should enroll in the school so the ratio becomes 4:3?





1
Expert's answer
2021-11-26T14:50:24-0500

1. Proportion:

2540 — 100 %

x — 60 %

"x = \frac{2540 \times 60}{100} = 1524"

The first price = 2540 + 1524 = php 4064

Proportion:

4064 — 100 %

y — 20 %

"y = \frac{4064 \times 20}{100} = 812.8"

The second price = 4064 - 812.8 = php 3251.2

Proportion:

3251.2 — 100 %

z — 10 %

"z = \frac{3251.1 \times 10}{100} = 325.12"

A. The final selling price of the golden casing = 3251.2 -325.12 = php 2926.08

2929.08-2540 = 386.08

Proportion:

2540 — 100%

386.08 — a

"a = \frac{386.08 \times 100}{2540} = 15.2 \; \%"

B. Mr. Llanto's markup rate based on cost is 15.2 %

2.

pckle relish: "0.4 \times 250 = php 100"

ham: "275.0 \times 0.5 = php 137.5"

cheese: "0.50 \times 200 = php 100"

bell pepper: "135 \times \frac{10}{1000} = php 1.35"

eggs: "7 \times 6 = php 42"

Total: php 380.85

1000 -380.85 = php 619.15

Her change will be php 619.15

3.

In school, there are 650 students. The ratio of the boys to the girls is 8:5. How many more girls should enroll in the school so the ratio becomes 4:3?

Real number of girls "= \frac{5}{8+5} \times 650 = 250"

New number of girls "= \frac{3}{4+3} \times 650 = 278"

Δ=278-250=28